Someone please help me solve this problem, I am going to have similar ones on a test. Find three consecutive even intergers such that three times the second minus the third is 12 more than the first.
Answer by Annabelle1(69)

let you integers be x, x+1, x+2
three times the second is 3(x+1) minus the third -(x+2)=12 more than the first (x+1)+12
so 3(x+1)-(x+2)=(x+1)+12
3x+3-x-2=x+1+12
2x+3=x+13
minus x from both sides
2x-x+3=x-x+13
x+3=13
minus 3 from both sides
x+3-3=13-3
x=10
the first integer is 10 so teh second and third (x+1) and (x+2) are 11 and 12